suppose that the market demand for aluminium in agiven country is given by the equation Qd =600-50p where P is the price of aluminium expressed in dollar per dilogram and the market supply for aluminium is given by the equation Q=-400 +50P and if the price of aluminium by10 % per kilogram from equilibrum price then


1
Expert's answer
2021-06-24T12:36:27-0400

Given the following;

Qd = 600 - 50P and Qs = -400 + 50P

At equilibrium, Qd = Qs

600 - 50P = -400 + 50p

600 + 400 = 50P + 50P

1000 = 100P

"\frac{1000}{100} = P"

P = 10

Equlibrium price = $10

Equilibrium Quantinty = 600 - (50 "\times" 10) or -400 + (50 "\times" 10)

= 600 - 500 or - 400 + 500 = 100units

After 10% increase, New Price = 10 + (10 "\times" 0.1) = $11


Effects on Quantity Demanded and Quantity supplied:


Qd = 600 - (50 "\times" 11) = 600 - 550 = 50units (decrease)


Qs = -400 + (50 "\times"11) = -400 + 550 = 150units (increase)